OpenStudy (anonymous):

If f(x) = 5 x^{5/2}, find f'( 5 ).

OpenStudy (anonymous):

do you know how to find f'(x) aka the first derivative of f(x)

OpenStudy (anonymous):

y, {f(x+h)-f(x)}/h

OpenStudy (anonymous):

or for \[x^n\] the derivative is \[n*x^{n-1}\]

OpenStudy (anonymous):

so\[f'(x) = 5( (5/2)x^{(\frac{ 5 }{ 2 } - 1)}) = (25/2)x^\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

thx a lot. i got.